T=2π√l\g,if l=50.0+\-0.2cm and the time for 15 oscillations are:24.2,24.2,24.7,24.5,24.2,24.2,24.4,24.5,24.6,24.7,24.5,24.6,24.5,24.3, determine the value of g and the standard error in g
1
Expert's answer
2020-12-29T15:32:32-0500

"l=(0.50\pm0.02)m"


For the period after statistical calculations


"T=(1.6\pm0.1)s"


"g=4\pi^2\frac{l}{T^2}=4\cdot3.14^2\cdot\frac{0.5}{1.6^2}=7.71(m/s^2)"


"\frac{\Delta g}{g}=\sqrt{(\frac{\Delta l}{l})^2+(2\frac{\Delta T}{T})^2}="


"=\sqrt{(\frac{0.02}{0.50})^2+(2\cdot\frac{0.1}{1.6})^2}=0.131"


"\Delta g=7.71\cdot0.131=1.01(m/s^2)"


"g=(7.7\pm1.0)m/s^2"